OpenStudy (anonymous):

Which of the following is a solution to tan2 x − 1 = 0?

OpenStudy (anonymous):

tan2(x)−1=0 tan2(x)=1 tan(x)=±1√ tan(x)=1√ or tan(x)=−1√ tan(x)=1 or tan(x)=−1 x=arctan(1) or x=arctan(−1) x=π4+πn or x=3π4+πn,x=−π4+πn or x=−3π4+πn These solutions can be combined into the following x=πn2−π4